A circular path around a lake has a diameter of 1,500 feet (ft). Pranotti jogs around the path 3 times for her morning run. How far does she run? Use 3.14 for and round your answer to the nearest foot.
Answer by RAY100(1637)

for 3 times around, let Distance = 3 * Circumference
.
d = 3 * 2 *pi * radius, radius = 1500 / 2 = 750
.
d = 3 * 2 * 3.14 * 750 = 14,130 ft